OpenStudy (anonymous):

line perpendicular to y=5x-4 and passes through the point (3,4)

OpenStudy (anonymous):

The slope of y=5x-4 is 5. We know this because its in the form y = mx + b We want to be perpendicular to this line. To get a perpendicular slope, take the negative reciprocal of 5. That just means flip it upside down and change the sign. The negative reciprocal of 5(which is really just 5/1) is -1/5. Now use the point slope equation y - y1 = m(x - x1) where (x1,y1) is any point on your line. For you x1 is 3 and y1 is 4.

OpenStudy (anonymous):

thank you so much, that helps a lot

OpenStudy (anonymous):

do you understand how to find the line ?

OpenStudy (anonymous):

BangkokGarrett explained it very well